The estimated life time of the sun on mainsequence is ~10^10 yrs. Determine the mainsequence life time of a star of 0.5 Mo.
1
Expert's answer
2019-06-03T09:50:21-0400

If you make a rough estimate of how much energy the sun emits, you’ll get something like:

"E=0.007 \times 0.1 \times M_{sun}c^2"

the rate at which the sun is giving off energy is equal "3.8 \times 10^{26} W" . And lifetime of the sun is

"t_{sun} = \frac{E}{3.8 \times 10^{26}} = \frac{0.007 \times 0.1 \times M_{sun}c^2}{3.8 \times 10^{26}}"

If mass of the star is "M_{star} = 0.5M_{sun}", then

"t_{star}=0.5t_{sun} = 5\times 10^9 yrs"